Sutton Pointers Corporation expects to begin operations on January 1, 2015; it will operate as a specialty sales company that sells laser pointers over the Internet. Sutton expects sales in January 2015 to total $280,000 and to increase 10 percent per month in February and March. All sales are on account. Sutton expects to collect 66 percent of accounts receivable in the month of sale, 21 percent in the month following the sale, and 13 percent in the second month following the sale.

Required
a. Prepare a sales budget for the first quarter of 2015.

     


b.

Determine the amount of sales revenue Sutton will report on the first 2015 quarterly pro forma income statement.

     

c.

Prepare a cash receipts schedule for the first quarter of 2015. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)

     

d.

Determine the amount of accounts receivable as of March 31, 2015. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)
Answer #1

a) Sales budget

January February March Total
Sales revenue 280000 280000*1.1 = 308000 308000*1.1 = 338800 926800

b) Sales revenue reported in income statement = 926800

c) Cash receipt budget

January February March Total
January Sales 184800 58800 36400 280000
February Sales 203280 64680 267960
March Sales 223608 223608
Total 184800 262080 324688 771568

d) Account receivable = (308000*13%+338800*34%) = $155232